OpenStudy (anonymous):

Solve for the x 3(-2x-1)=2(x+5) -13/8 -3/4 13/8 3/4

hartnn (hartnn):

distribute using a(b+c) = ab+ac what u get?

OpenStudy (anonymous):

um yea i suck at math so idk

hartnn (hartnn):

3*(-2x)=?

OpenStudy (anonymous):

60

hartnn (hartnn):

how?

OpenStudy (anonymous):

calc. haha

hartnn (hartnn):

3(-2x-1)=2(x+5) -6x-3 = 2x +10 -3-10 = 2x+6x 8x = -13 x= -13/8 ask if there is doubt in any step.

OpenStudy (anonymous):

oooo ok i get it thanx

hartnn (hartnn):

welcome ^_^